OpenStudy (anonymous):

What is the solution of log(2x + 6) 8 = 3?

OpenStudy (anonymous):

Is it x=-2?

OpenStudy (solomonzelman):

(2x+6) is the base ?

OpenStudy (anonymous):

Yes :)

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \log_{(2x+6)}8=3 }\] CLOSE ATTENTION ! \[\huge\color{blue}{ \log_{(2x+6)}8=3 }\huge\color{red}{ \times \log_{(2x+6)}(2x+6) } \] \[\huge\color{blue}{ \log_{(2x+6)}8=}\huge\color{blue}{ \log_{(2x+6)}(2x+6)^3 } \] \[\huge\color{blue}{ 8=(2x+6)^3 } \] see what I did and can do it from there ?

OpenStudy (anonymous):

I think so! Would it be...

OpenStudy (solomonzelman):

Can you (to make sure) explain to me what I did in red ?

OpenStudy (anonymous):

I think so! Would it be...\[2^3 = (2x+6)^3\]

OpenStudy (solomonzelman):

Yep, good step!

OpenStudy (anonymous):

I'm honestly not sure.. did you set it up so both sides equal each other?

OpenStudy (solomonzelman):

yes, what I did in red is based on a rule \[\log_xx=1\]

OpenStudy (anonymous):

What is the rule called?

OpenStudy (solomonzelman):

I don't remember.